2x-8y=16
-x+4y=-8

Answer by lpilon(9)

2x - 8y = 16
-x + 4y = -8
_____________
x - 2y = 8
x = 2y + 8
So, substitute the value of x into one of the equations:
2(2y + 8) - 8y = 16
4y + 16 - 8y = 16
Solve for y
-4y = 16 - 16
-4y = 0
y = 0
In the other equation using the value of x
-(2y + 8)+ 4y = -8
-2y - 8 + 4y = -8
Solve for y
2y = -8 + 8
y = 0